### v1

@@ 4.2 Profiling Harness
We develop an end-to-end, cross-platform profiling harness for inference workloads, ensuring reproducible results and support for new models, tasks, and hardware backends. It consists of three core components: distributed multi-GPU inference, response evaluation, and system-level telemetry collection. The harness supports three hardware backends: NVIDIA, macOS (Apple Silicon), and AMD. Given a dataset, model, and hardware backend, it orchestrates inference over all input queries, evaluates outputs (via exact match or LLM-as-a-judge), and records detailed telemetry—latency, throughput, time-to-first-token (TTFT), energy consumption, and more (see Table 8). Telemetry is collected via vendor APIs, synchronized at nanosecond resolution, and normalized (watts, joules, megabytes). For energy measurements, we follow standard practices (Samsi et al., 2023a; Fernandez et al., 2025; Wilkins et al., 2024): querying NVML’s energy counter for NVIDIA GPUs and using powermetrics with power integration over time for Apple Silicon. We differ in two key aspects: (1) we sample at 50ms intervals for higher temporal resolution (compared to 100ms (Samsi et al., 2023a) or 15 seconds (Fernandez et al., 2025)), and (2) for multi-GPU configurations, we aggregate energy counters from each GPU individually rather than extrapolating from a single device (Samsi et al., 2023a). On NVIDIA systems, we use NVML to retrieve per-device power, energy, memory usage, and temperature. On macOS systems, we extract GPU power data from powermetrics and compute energy via numerical integration. On AMD systems, we query ROCm SMI for power, temperature, and VRAM usage, again integrating power over time. Full details provided in App. B.1.
@@ 5 Intelligence Efficiency Study
Figure 2: Local Models Rival Cloud Models Across Diverse Benchmarks: Individual model performance scales with size, ranging from 28.8–72.5% for Qwen3-4B to 66.5–89.5% for Gemini 2.5 Pro. Local routing (best local LM per query) achieves 97.8%, 88.3%, 77.0%, and 92.4% on Wildchat, NaturalReasoning, SuperGPQA, and MMLU Pro respectively, surpassing cloud routing (100%, 82.9%, 66.5%, 87.4%) on three of four benchmarks.
@@ 5.2 How Intelligence Efficient is Local Inference?
$7.92\times 10^{-4}$ $1.80\times 10^{-3}$ $4.18\times 10^{-3}$
@@
$1.40\times 10^{-3}$ $1.63\times 10^{-3}$ $1.69\times 10^{-3}$ $1.97\times 10^{-3}$
@@
$1.95\times 10^{-3}$ $2.27\times 10^{-3}$ $2.35\times 10^{-3}$ $2.75\times 10^{-3}$
@@
$3.80\times 10^{-5}$ $3.51\times 10^{-5}$ $4.38\times 10^{-5}$ $4.23\times 10^{-5}$
@@
$8.71\times 10^{-5}$ $5.91\times 10^{-5}$ $7.34\times 10^{-5}$ $6.78\times 10^{-5}$
@@ 8 Acknowledgements
We thank Simran Arora, Bradley Brown, Mayee Chen, Owen Dugan, Neel Guha, Simon Guo, Jordan Juravsky, Jerry Liu, Benjamin Spector, and Shayan Talaei for their constructive feedback during the composition of the paper. We would also like to thank our collaborators at the Stanford Artificial Intelligence Laboratory (SAIL) and TogetherAI.

### v2

@@ 4.2 Profiling Harness
We develop an end-to-end, cross-platform profiling harness for inference workloads that ensures reproducible results and easily accommodates new models, tasks, and hardware backends. It comprises three components—distributed multi-GPU inference, response evaluation, and system-level telemetry collection—and currently supports NVIDIA, macOS (Apple Silicon), and AMD systems. Given a dataset, model, and backend, the harness orchestrates inference over all input queries, evaluates outputs (via exact match or LLM-as-a-judge), and records detailed telemetry—latency, throughput, time-to-first-token (TTFT), energy consumption, and more (Table 8). Telemetry is collected via vendor APIs, synchronized at nanosecond resolution, and normalized to common units (watts, joules, megabytes). For energy measurements, we follow standard practices (Samsi et al., 2023a; Fernandez et al., 2025; Wilkins et al., 2024). On NVIDIA systems we query NVML for per-device power, energy, memory usage, and temperature; on macOS systems we extract GPU power data from powermetrics; and on AMD systems we query ROCm SMI for power, temperature, and VRAM usage. In all cases, we compute energy via numerical integration over time and sample at 50 ms intervals, providing higher temporal resolution than prior work (100 ms (Samsi et al., 2023a) or 15 s (Fernandez et al., 2025)). For multi-GPU configurations, we aggregate energy from each GPU individually rather than extrapolating from a single device (Samsi et al., 2023a). However, software-based power measurements can introduce inaccuracies of 10–15%, with variations distributed across different hardware components due to architectural differences in workloads between CPUs, GPUs, and NPUs (Yang et al., 2023). Even hardware wattage meters may fall short for milliwatt-level precision, though our approach aligns with established practices and provides consistent relative comparisons across configurations. Full implementation details are provided in App. B.1.
@@ 5 Intelligence Efficiency Study
Figure 2: Local Models Rival Cloud Models Across Diverse Benchmarks: Individual model performance scales with size, ranging from 31.5–69.4% for IBM Granite4-H-Small, 30.0–83.6% for Gemma3-12B, 51.5–80.4% for GPT-OSS-120B, and 66.5–89.5% for Gemini 2.5 Pro. Local routing (best local LM per query) achieves 97.8%, 88.3%, 77.0%, and 92.4% on Wildchat, NaturalReasoning, SuperGPQA, and MMLU Pro respectively, surpassing cloud routing (100%, 82.9%, 66.5%, 87.4%) on three of four benchmarks.
@@ 5.2 How Intelligence Efficient is Local Inference?
$(7.92\pm 0.32)$
$\times 10^{-4}$
$(1.80\pm 0.21)$
$\times 10^{-3}$
$(4.18\pm 0.53)$
$\times 10^{-3}$
@@
$(1.40\pm 0.38)$
$\times 10^{-3}$
$(1.63\pm 0.20)$
$\times 10^{-3}$
$(1.69\pm 0.31)$
$\times 10^{-3}$
$(1.97\pm 0.24)$
$\times 10^{-3}$
@@
$(1.95\pm 0.14)$
$\times 10^{-3}$
$(2.27\pm 0.18)$
$\times 10^{-3}$
$(2.35\pm 0.09)$
$\times 10^{-3}$
$(2.75\pm 0.14)$
$\times 10^{-3}$
@@
$(3.80\pm 0.40)$
$\times 10^{-5}$
$(3.51\pm 0.38)$
$\times 10^{-5}$
$(4.38\pm 0.31)$
$\times 10^{-5}$
$(4.23\pm 0.45)$
$\times 10^{-5}$
@@
$(8.71\pm 0.60)$
$\times 10^{-5}$
$(5.91\pm 0.42)$
$\times 10^{-5}$
$(7.34\pm 0.51)$
$\times 10^{-5}$
$(6.78\pm 0.47)$
$\times 10^{-5}$
@@ 8 Acknowledgements
We thank Simran Arora, Bradley Brown, Mayee Chen, Owen Dugan, Sabri Eyuboglu, Neel Guha, Simon Guo, Jordan Juravsky, Jerry Liu, Benjamin Spector, Shayan Talaei, and Dan Zhang for their constructive feedback during the composition of the paper. We would also like to thank our collaborators at the Stanford Artificial Intelligence Laboratory (SAIL) and TogetherAI.
@@ 94. Technical precision
To ensure measurement precision and account for variance in inference-time behavior, we execute each query 10 times and aggregate power measurements across runs. For each query, we compute the mean power draw (watts) and mean energy consumption (joules) per query by averaging across these 10 independent executions. This repeated sampling approach reduces measurement noise and provides robust estimates of per-query resource consumption that account for system-level variability in accelerator utilization, thermal conditions, and memory allocation patterns.